Please I need your help to explain the answers for the following question

Wyatt Oil, an all-equity financed firm, has just reported EPS of $4.00 per share. Despite an economic downturn, Wyatt is confident regarding its current investment opportunities, but due to the current financial crisis, Wyatt does not wish to fund these investments externally. Wyatt's board has therefore decided to suspend its stock repurchase plan and cut its dividend to $1 per share (from its current level of $2 per share) and retain these funds instead. The firm just paid its current dividend of $1.00 per share and expects to keep its dividend at $1 per share next year as well. In subsequent years, it expects its growth opportunities to slow, and it will still be able to fund its growth internally with a target 40% dividend payout ratio, and reinitiating its stock repurchase plan for a total payout rate of 60%. All dividends and repurchases occur at the end of each year.
Wyatt's existing operations are expected to generate the current level of earnings per share in the future. Assume that the return on new investments is 16% and that reinvestments will account for all future earnings growth. Wyatt's current equity cost of capital is 12%.

13) Wyatt's expected EPS in two years is closest to:
A) $4.48
B) $4.64
C) $5.04
D) $5.38
Answer: C

14) Wyatt's current stock price is closest to:
A) $51.23
B) $54.00
C) $49.11
D) $61.38
Answer: C
Answer #1

Ans:-

13)Year 1:-

Retention ratio =(Net income - Dividend) /Net Income

Retention ratio = (4 - 1)/4 = 0.75

Growth rate = Retention ratio × ROE

Growth rate = 0.75 ×16% = 0.12 or 12%

ROE =return on new investments

EPS = $4.00(1 + 12%)

EPS = $4.00(1.12)

EPS =$4.48

Year 2:- Retention Ratio =(Net income - Dividend) /Net Income

Retention ratio =(4.48 - 1)/4.48 =0.776786

Growth rate =Retention ratio ×ROE

Growth rate =0.776786 ×16% =0.1243 or 12.43%

EPS = $4.48(1 + 12.43%)

EPS =$4. 48(1.1243)

EPS =$5.04

Thus the option C is Right Answer.

14)Total payout ratio (TPR) =60%

Retention rate (RR) = 1- TPR = 1 – 60% = 40%

Growth = RR ×ROI =40% × 16% =6.4%

Dividend =5.04 × 60%= 3.024

Current Stock Price =2.750 /(0.12 - 0.064)

Current Stock Price =$49.11